OpenStudy (anonymous):

What is the factored form of the following expression? w2 + 18w + 77 A. (w – 7)(w + 11) B. (w – 7)(w – 11) C. (w + 7)(w + 11) D. (w + 1)(w + 77)

Parth (parthkohli):

w2 + 11w + 7w + 77 = w(w + 11) + 7(w + 11) = (w + 7)(w + 11)